An object moves 60m in East and than 80m in North find the distance travelled and the magnitude of displacement?
1
Expert's answer
2018-04-04T15:41:10-0400
distance = 60 + 80 = 140 m
|displacement| = sqrt(60^2 + 80^2) = 100 m